# Weixiang Zhao1, Yulin Hu1   Jiahe Guo1footnotemark:    Xingyu Sui1footnotemark:   , Tongtong Wu2, Yang Deng3 Yanyan Zhao1, Bing Qin1, Wanxiang Che1, Ting Liu1 1

### v1

@@ 6 Conclusion
In this paper, we introduce Lens, a novel method designed for the effective, efficient and comprehensive multilingual enhancement of large language models (LLMs). Lens first decouple the multilingual hidden spaces of the backbone into two orthogonal components: a language-agnostic subspace and a language-specific subspace. Then taking well-established representations of the central language as a pivot, representations of target languages are pulled closer and pushed away from them in language-agnostic subspace and language-specific subspace, respectively. Experimental results on 3 representative cutting-edge LLMs demonstrate that Lens outperforms baseline methods with much lower training costs, underscoring its efficacy, efficiency and scalability.
Appendix A Limitation and Future Work
@@
Appendix B Probing for Language Subspace
@@
In: languages’ mean embeddings $\bm{M}$, rank of subspace $r$
Out: language-agnostic subspace $\bm{M}_{a}$, language-specific subspace $\bm{M}_{s}$, coordinates $\bm{\Gamma}$
/* 1) Approximate $\bm{M}$ in low rank */
1 $\bm{M}_{a}^{\prime}\leftarrow\frac{1}{d}\bm{M}\bm{\mathbbm{1}}$;
2 $\bm{M}_{s}^{\prime},\text{\_},\bm{\Gamma}^{\prime}\leftarrow\text{Top-}r\text{SVD}\left(\bm{M}-\bm{M}_{a}^{\prime}\bm{\mathbbm{1}}^{\top}\right)$;
3 $\bm{M}^{\prime}\leftarrow\bm{M}_{a}^{\prime}\bm{\mathbbm{1}}^{\top}+\bm{M}_{s}^{\prime}{\bm{\Gamma}^{\prime}}^{\top}$;
@@ 1 $\bm{M}_{a}^{\prime}\leftarrow\frac{1}{d}\bm{M}\bm{\mathbbm{1}}$;
4 $\bm{M}_{a}\leftarrow\frac{1}{\|{\bm{M}^{\prime}}^{+}\bm{\mathbbm{1}}\|^{2}}{\bm{M}^{\prime}}^{+}\bm{\mathbbm{1}}$;
$\bm{M}_{s},\text{\_},\bm{\Gamma}\leftarrow\text{Top-}r\text{ SVD}\left(\bm{M}^{\prime}-\bm{M}_{a}\bm{\mathbbm{1}}^{\top}\right)$
@@ Appendix C Multilingual Benchmarks
  • XWinograd (Muennighoff et al., 2023): [^7] https://huggingface.co/datasets/Muennighoff/xwinograd A benchmark to evaluate the ability of machine learning models to transfer commonsense reasoning across languages. The dataset is the translation of the English Winograd Schema datasets and it adds 488 Chinese schemas from CLUEWSC2020 , totaling 6 languages. Formulated as a fill-in-a-blank task with binary options, the goal is to choose the right option for a given sentence which requires commonsense reasoning. In our experimental setup, this benchmark covers English (En), Chinese (Zh) and Japanese (Jp).
  • XStoryCloze (Lin et al., 2022): [^8] https://huggingface.co/datasets/juletxara/xstory_cloze A benchmark to evaluate the ability of machine learning models to transfer commonsense reasoning across languages. The dataset consists of the professionally translated version of the English StoryCloze dataset (Spring 2016 version) to 10 non-English languages. The dataset is challenging and is designed to evaluate story understanding, story generation, and script learning. In our experimental setup, this benchmark covers English (En), Chinese (Zh), Arabic (Ar) and Swahili (Sw).
  • M-MMLU (Hendrycks et al., 2021; Lai et al., 2023): [^9] https://huggingface.co/datasets/alexandrainst/m_mmlu A benchmark to evaluate the ability of machine learning models to transfer commonsense reasoning across languages. The datasets is a machine translated version of the MMLU dataset by GPT-3.5-turbo and covers 34 languages. This is a massive multitask test consisting of multiple-choice questions from various branches of knowledge. To attain high accuracy on this test, models must possess extensive world knowledge and problem solving ability. In our experimental setup, this benchmark covers English (En), Chinese (Zh), Arabic (Ar), Korean (Ko), and Swahili (Sw).
@@
Phi-3.5-mini-Instruct 29 0.2 0.2 0.2 0.2 0.2 0.2
For multilingual generation:
Appendix D Implementation Details
@@
More detailed hyper-parameter settings for bilingual and multilingual enhancement across different backbones are listed in Table 2 and Table 3, respectively.
Further, we carefully evaluate the official implementations of all baselines, in order to make the comparison as fair as possible. We strictly follow the hyper-parameter settings in their original code. If this could not reach the expected performance, we carry out the hyper-parameter search of the learning rate and batchsize.
Table 4: Detailed results on the multilingual understanding and multilingual generation benchmarks with Phi-3.5-mini-Instruct backbone under the bilingual setting (English and Chinese). Accuracy serves as the evaluation metric for multilingual understanding, while GPT-4o ratings (on a scale of 1 to 10) are provided for MT-Bench. The values in parentheses represent language fidelity. Results highlighted in green indicate an improvement or performance comparable to the original backbone, while those highlighted in red signal a decline in performance relative to the original backbone.
@@
Lens (Ours) - 81.60 75.80 67.66 95.40 89.40 71.70 47.40 80.97 71.51 6.44 (100%) 5.16 (92.50%)
@@
Lens xSFT xSFT-Full SDRRL QAlign V1 V2 V3
Training time 2m08s 5m33s 192m35s 11m30s 12m03s - - -
Trainable parameters rate 5.43% 100.00% 100.00% 100.00% 100.00% 13.08% 13.08% -
Instruction data 0.8K 0.8K 111.5K 4K 0.8K 5M 5M 5K
Pre-training data - - - - - 120G - -
Figure 7: Results on the multilingual understanding benchmark with LLaMA-3-8B-Instruct backbone under the multilingual setting. We report the average performance of each language on the corresponding benchmarks.
@@
Appendix E Additional Experimental Results
@@
We report the multilingual understanding performance of LLaMA-3-8B-Instruct in Figure 7. Experimental results of the comparison between Lens and baseline methods on Phi-3.5-mini-Instruct under bilingual and multilingual setups are shown in Table 4 and Figure 9, respectively. And the multilingual enhancement results for LLaMA-3.1-8B-Instruct are displayed in Figure 8.
@@
The results demonstrate that our Lens is still capable of achieving the comprehensive multilingual enhancement. Similarly, Lens continues to improve the model’s multilingual generation capability, enhancing the quality of the model’s responses in specific languages. However, the improvement in language fidelity is more pronounced in the English-centric backbone than in the multilingual backbone, which the latter one undergoes more extensive multilingual alignment training. Notably, while the baseline method considerably decreases the language fidelity of the multilingual backbone, Lens has minimal impact on it. These extensive experimental results demonstrate that Lens can serve as an effective, efficient, and scalable multilingual enhancement solution. We hope that our method can provide inspiration for future work to seek multilingual supervision more from the LLM itself rather than heavily relying on external dataset.

### v2

@@ 6 Conclusion
In this paper, we introduce Lens, a novel method designed for the effective, efficient and comprehensive multilingual enhancement of large language models (LLMs). Lens first decouples the multilingual hidden spaces of the backbone into two orthogonal components: a language-agnostic subspace and a language-specific subspace. Then taking well-established representations of the central language as a pivot, representations of target languages are pulled closer and pushed away from them in language-agnostic subspace and language-specific subspace, respectively. Experimental results on 3 representative cutting-edge LLMs demonstrate that Lens outperforms baseline methods with much lower training costs, underscoring its efficacy, efficiency and scalability.
Limitation and Future Work
@@
Ethical Considerations
This work is conducted solely for academic research purposes, aiming to enhance the multilingual capabilities of large language models in a responsible and interpretable manner. Our methods focus on improving cross-lingual alignment and representation learning without introducing biases or harmful content. We do not intend for this research to be used in applications that could lead to misinformation, discrimination, or unethical manipulation of language models. Additionally, we adhere to ethical AI principles, ensuring that our approach respects linguistic diversity and maintains fairness across different languages.
Appendix A Probing for Language Subspace
@@
In: languages’ mean embeddings $\boldsymbol{M}$, rank of subspace $r$
Out: language-agnostic subspace $\boldsymbol{M}_{a}$, language-specific subspace $\boldsymbol{M}_{s}$, coordinates $\boldsymbol{\Gamma}$
/* 1) Approximate $\boldsymbol{M}$ in low rank */
1 $\boldsymbol{M}_{a}^{\prime}\leftarrow\frac{1}{d}\boldsymbol{M}\boldsymbol{\mathbbm{1}}$;
2 $\boldsymbol{M}_{s}^{\prime},\text{\_},\boldsymbol{\Gamma}^{\prime}\leftarrow\text{Top-}r\text{ SVD}\left(\boldsymbol{M}-\boldsymbol{M}_{a}^{\prime}\boldsymbol{\mathbbm{1}}^{\top}\right)$;
3 $\boldsymbol{M}^{\prime}\leftarrow\boldsymbol{M}_{a}^{\prime}\boldsymbol{\mathbbm{1}}^{\top}+\boldsymbol{M}_{s}^{\prime}{\boldsymbol{\Gamma}^{\prime}}^{\top}$;
@@ 1 $\boldsymbol{M}_{a}^{\prime}\leftarrow\frac{1}{d}\boldsymbol{M}\boldsymbol{\mathbbm{1}}$;
4 $\boldsymbol{M}_{a}\leftarrow\frac{1}{\|{\boldsymbol{M}^{\prime}}^{+}\boldsymbol{\mathbbm{1}}\|^{2}}{\boldsymbol{M}^{\prime}}^{+}\boldsymbol{\mathbbm{1}}$;
$\boldsymbol{M}_{s},\text{\_},\boldsymbol{\Gamma}\leftarrow\text{Top-}r\text{ SVD}\left(\boldsymbol{M}^{\prime}-\boldsymbol{M}_{a}\boldsymbol{\mathbbm{1}}^{\top}\right)$
@@ Appendix C Multilingual Benchmarks
  • XWinograd (Muennighoff et al., 2023): [^2] https://huggingface.co/datasets/Muennighoff/xwinograd A well-established tool for evaluating coreference resolution (CoR) and commonsense reasoning (CSR) capabilities of computational models. The dataset is the translation of the English Winograd Schema datasets and it adds 488 Chinese schemas from CLUEWSC2020 (Xu et al., 2020), totaling 6 languages. Formulated as a fill-in-a-blank task with binary options, the goal is to choose the right option for a given sentence which requires commonsense reasoning. In our experimental setup, this benchmark covers English (En), Chinese (Zh) and Japanese (Jp).
  • XStoryCloze (Lin et al., 2022): [^3] https://huggingface.co/datasets/juletxara/xstory_cloze A commonsense reasoning framework for evaluating story understanding, story generation, and script learning. The dataset consists of the professionally translated version of the English StoryCloze dataset (Spring 2016 version) to 10 non-English languages. The dataset is challenging and is designed to evaluate story understanding, story generation, and script learning. In our experimental setup, this benchmark covers English (En), Chinese (Zh), Arabic (Ar) and Swahili (Sw).
  • M-MMLU (Hendrycks et al., 2021; Lai et al., 2023): [^4] https://huggingface.co/datasets/alexandrainst/m_mmlu A benchmark designed to measure knowledge acquired during pretraining by evaluating models exclusively in zero-shot and few-shot settings. The datasets is a machine translated version of the MMLU dataset by GPT-3.5-turbo and covers 34 languages. This is a massive multitask test consisting of multiple-choice questions from various branches of knowledge. To attain high accuracy on this test, models must possess extensive world knowledge and problem solving ability. In our experimental setup, this benchmark covers English (En), Chinese (Zh), Arabic (Ar), Korean (Ko), and Swahili (Sw).
For multilingual generation:
Appendix D Baseline Methods
D.1 Multilingual Enhancement Baselines
For comparison, we consider the following baseline methods that enhance LLMs’ multilingual capabilities using multilingual instruction fine-tuning technique: (1) xSFT & xSFT-Full (Ouyang et al., 2022): xSFT performs multilingual instruction fine-tuning using the same data volume as our Lens. In contrast, xSFT-Full utilizes the full dataset for each target language from the Aya Collection and Bactrian-X. (2) QAlign (Zhu et al., 2024a): It explores the benefits of question alignment, where the model is trained to translate inputs into English by finetuning on X-English parallel question data. (3) SDRRL (Zhang et al., 2024c): It is based on self-distillation from resource-rich languages that effectively improve multilingual performance by leveraging self-distillated data. (4) CLA (Li et al., 2024): It aligns the internal sentence representations across different languages via multilingual contrastive learning and aligns outputs by following cross-lingual instructions in the target language.
D.2 Open-Source Multilingual-Enhanced LLMs
The Chinese-LLaMA-3 series, which builds on LLaMA-3 series to enhance Chinese capabilities and includes three different versions:
@@ D.2 Open-Source Multilingual-Enhanced LLMs
Phi-3.5-mini-Instruct 29 0.2 0.2 0.2 0.2 0.15 0.15
Appendix E Implementation Details
@@
More detailed hyper-parameter settings for bilingual and multilingual enhancement across different backbones are listed in Table 3 and Table 4.
Further, we carefully evaluate the official implementations of all baselines, in order to make the comparison as fair as possible. All baselines are trained using the same language pairs as those in Lens. For SDRRL and CLA, we adhere to their respective configurations to reconstruct the training data from these language pairs. We strictly follow the hyper-parameter settings in their original code. If this could not reach the expected performance, we carry out the hyper-parameter search of the learning rate and batch size.
@@
CLA - 80.40 76.50 66.50 95.70 89.40 71.70 47.10 81.30 70.85 6.08 (100%) 4.26 (90.00%)
Lens (Ours) - 82.60 75.80 68.73 95.60 90.60 71.80 47.40 81.07 72.33 6.44 (100%) 5.16 (92.50%)
Table 5: Detailed results on the multilingual understanding and multilingual generation benchmarks with Phi-3.5-mini-Instruct backbone under the bilingual setting (English and Chinese). Accuracy serves as the evaluation metric for multilingual understanding, while GPT-4o ratings (on a scale of 1 to 10) are provided for MT-Bench. The values in parentheses represent language fidelity. Results highlighted in green indicate an improvement or performance comparable to the original backbone, while those highlighted in red signal a decline in performance relative to the original backbone. The best and second-best results in our method and baselines are in bold and underlined.
Lens xSFT xSFT-Full SDRRL QAlign CLA V1 V2 V3
Training time 2m08s 5m33s 192m35s 11m30s 12m03s 12m33s - - -
Trainable parameters rate 5.43% 100.00% 100.00% 100.00% 100.00% 100.00% 13.08% 13.08% -
Instruction data 1K 1K 111.5K 4K 1K 1K 5M 5M 5K
Pre-training data - - - - - - 120G - -
@@
Figure 6: Results on the multilingual understanding and generation benchmarks with LLaMA-3-8B-Instruct backbone under the multilingual setting.
@@
Appendix F Additional Experimental Results
F.1 Results under Multilingual Settings
@@
Results of the comparison between Lens and baseline methods on Phi-3.5-mini-Instruct under bilingual and multilingual setups are shown in Table 5 and Figure 8, respectively. We report the results of multilingual enhancement settings on LLaMA-3-8B-Instruct and LLaMA-3.1-8B-Instruct are displayed in Figure 6 and 7, respectively.
@@
The results demonstrate that our Lens is still capable of achieving the comprehensive multilingual enhancement. Similarly, Lens continues to improve the model’s multilingual generation capability, enhancing the quality of the model’s responses in specific languages. However, the improvement in language fidelity is more pronounced in the LLaMA-3-8B-Instruct backbone than in the multilingual backbones (LLaMA-3.1-8B-Instruct and Phi-3.5-mini-Instruct), which the latter ones undergo more extensive multilingual alignment training. Notably, while the baseline method considerably decreases the language fidelity of the multilingual backbones, Lens has minimal impact on it. These extensive experimental results demonstrate that Lens can serve as an effective, efficient, and scalable multilingual enhancement solution. We hope that our method can provide inspiration for future work to seek multilingual supervision more from the LLM itself rather than heavily relying on external dataset.
We also note that xSFT-Full exhibits uneven performance gains, especially in Swahili (Sw). We believe this may be attributed to the uneven quality of the training data. Specifically, the Bactrian-X dataset used for training derives its input from Google Translate and its output from GPT-3.5-turbo, meaning the dataset quality depends heavily on these two sources. As a result, inconsistencies in translation and generation quality can introduce noise, leading to uneven performance gains from data-driven post-training approaches like xSFT-Full. This highlights one of the key limitations of the current data-driven paradigms.
In contrast, Lens seeks supervision signals internally from the backbone itself, bypassing the need for extensive reliance on potentially noisy external datasets. This intrinsic approach allows Lens to achieve consistent improvements over the backbone model across a wide range of languages, demonstrating better scalability and robustness. We have also demonstrated this phenomenon in our experiments, showcasing Lens’s broader applicability. In our future work, we propose combining the Lens training paradigm with advancements in data selection and filtering methods. We believe this hybrid approach holds great potential for further enhancing multilingual performance.
F.2 Comparison with LoRA-Based SFT
In Figures 6, 7 and 8, We also perform comparison with LoRA-based (Hu et al., 2022) SFT as recent work suggests that LoRA-based SFT can effectively prevent catastrophic forgetting (Huang et al., 2024). Based on our experimental results, we derived the following key conclusions:
For preserving the central language’s capabilities, incorporating LoRA-based SFT is indeed more effective at preventing catastrophic forgetting than its full-parameter counterpart. However, it primarily protects multilingual understanding (MU) tasks while multilingual generation (MG) capabilities are also significantly affected.
For target language enhancement, LoRA-based SFT methods also show a trend for improving MU tasks more over MG tasks.
By contrast, our proposed Lens achieves a more comprehensive performance, simultaneously enhancing understanding and generation for target languages while maintaining both the understanding and generation capabilities of the central language across different base models.
F.3 Broader Insights and Connection to Superficial Alignment Hypothesis
Our experimental findings suggest that eliciting language-specific representations benefits multilingual capability more than aligning language-agnostic representations, which may also lend support to the superficial alignment hypothesis (Zhou et al., 2023; Lin et al., 2024a; Yan et al., 2024). It posits that LLMs acquire their core knowledge and abilities during pretraining, while post-alignment training primarily guides the model towards a desirable subdistribution of formats to use when prompted. In the multilingual settings, this is specifically manifested in:
(1) Despite the imbalance distribution of training data in pretraining resources for different languages, the majority of language-agnostic knowledge is already well-comprehended and aligned during pretraining, especially for current LLMs exposed to super-large-scale pretraining corpora (e.g., over 15T tokens for LLaMA-3).
(2) Current post-alignment training, which disproportionately focuses on English data, limits other languages to a subdistribution aligned with English-specific formats.
Thus, further aligning multilingual representations may have less impact compared to stimulating language-specific expressiveness in the target languages, but both mechanisms contribute to performance improvement in our method, with separation playing a more significant role.
F.4 Results on Multilingual Machine Translation
We carry out evaluations on the FLORES-101 dataset (Goyal et al., 2022). Specifically, we assess the bidirectional translation performance between the target language and English, reporting scores using the COMET metric with the WMT22-comet-da model (Rei et al., 2022). The choice of target language and training settings align with the experimental setup used for multilingual enhancement on LLaMA-3-8B-Instruct.
The experimental results in Table 2 demonstrate that Lens still effectively enhances the multilingual machine translation performance, further validating its robustness across diverse multilingual tasks. Additionally, methods like QAlign, SDRRL and CLA, which heavily rely on translation-based training for multilingual alignment, fall short in acquiring deep linguistic understanding for each language. Thus, their ignorance of language-specific modeling hinders their efficacy in multilingual machine translation that require fine-grained linguistic nuances necessary for accurate translations.
F.5 Impact of Varying the Number of Manipulated Layers
Recent studies on the interpretability of LLMs has sought to reveal the mechanisms underlying their multilingual capabilities (Zhao et al., 2024c; Zhong et al., 2024). A growing consensus suggests that language-specific parameters or neurons are primarily concentrated in the upper layers of these models, while the middle layers tend to process inputs from various languages using a shared and language-agnostic mechanism (Chen et al., 2023a; Wendler et al., 2024; Tang et al., 2024; Kojima et al., 2024; Zhang et al., 2024d). Drawing inspiration from this, our main experiments focus on performing updates solely within the upper layers of the backbone, resulting in a notable improvement in multilingual performance. In Figure LABEL:subfig:layer_analysis, we explore the effect of increasing the number of layers involved on the model’s multilingual enhancement. The horizontal axis represents the starting point of the layers where manipulation is applied, with the default endpoint being the final layer. This experiment is performed under the bilingual enhancement with LLaMA-3-8B-Instruct.
It is important to note that modifying only the final layer does not significantly improve either multilingual understanding or generation. This is because language-specific information is not sufficiently enhanced, causing the model to suffer from off-target issues and struggle to represent specific languages accurately. The lack of improvement in multilingual understanding aligns with the findings in §4.1, which highlight the critical role of supervision provided by the Push loss ($\mathcal{L}_{2}$).
Lens further validates the conclusions of existing works on LLM interpretability and applies these findings to multilingual enhancement.
MU MG
Es Fr De Es Fr De
LLaMA-3 53.50 52.50 56.50 5.88 5.27 4.56
xSFT 50.80 49.90 56.00 4.36 4.47 4.00
xSFT-Full 50.90 48.90 51.50 4.60 4.42 4.33
SDRRL 48.90 47.80 50.30 2.74 3.06 2.54
QAlign 48.50 46.60 51.30 2.88 2.91 2.31
Lens 53.70 52.10 57.10 5.90 5.63 4.90
Table 7: Detailed results on the multilingual understanding and multilingual generation benchmarks with LLaMA-3-8B-Instruct backbone for Spanish (Es), French (Fr), and German (De). Accuracy serves as the evaluation metric for multilingual understanding, while GPT-4o ratings (on a scale of 1 to 10) are provided for MT-Bench. The values in parentheses represent language fidelity. The best and second-best results in our method and baselines are in bold and underlined.
F.6 Impact of Language Proximity
Here are our supplemented experimental results based on LLaMA-3-8B-Instruct to further highlight the generalizability of Lens, where these 3 languages, Spanish (Es), French (Fr), and German (De), are typologically and scripturally closer to English. For Multilingual Understanding (MU) evaluation, we adopt M-MMLU dataset which covers all 4 languages En, Es, Fr and De. And Multilingual Generation (MG) evaluation is performed on MT-Bench. Results are shown in Table 7.
The horizontal comparison in Table 7 and Figure 6 reveals that it demonstrates effectiveness across languages from different typological families, further highlighting its generalizability and adaptability to diverse linguistic characteristics.
Appendix G Additional Related Works
Here we provide additional discussion on the theoretical foundation of language-agnostic and language-specific subspaces, dividing it into two aspects: linguistic theory and LLM interpretability.
Linguistic Theory
From a linguistic standpoint, the idea of separating representations into language-agnostic and language-specific spaces is grounded in established theories of language universals and typology. Language-agnostic features align with universal linguistic structures, such as shared syntactic patterns or semantic primitives (Greenberg, 1963; Comrie, 1989), while language-specific features capture unique aspects like phonology, morphology, or syntax (Croft, 2002; Cotterell et al., 2016). These distinctions have also been studied in computational linguistics, such as in multilingual embeddings (Artetxe et al., 2018) and cross-lingual representation learning (Ruder et al., 2019), supporting the conceptual basis in Lens.
LLM Interpretability
Recent interpretability studies have provided compelling evidence that LLMs internally encode language-agnostic and language-specific subspaces. For example, specific neurons or groups of neurons have been identified as responsible for mapping multilingual input representations into either a shared language-agnostic space (Chen et al., 2023a; Starace et al., 2023; Wang et al., 2024b; Chen et al., 2024; Wendler et al., 2024) that different languages share the common knowledge or distinct language-specific spaces (Tang et al., 2024; Kojima et al., 2024; Zhang et al., 2024d) that are crucial for the accurate expression for specific languages. These findings support our assumption that LLMs naturally exhibit such separable structures, and our work leverages this inductive bias to improve multilingual performance.
Building upon such two theoretical foundations, particularly from linguistic theory, most previous works regarding multilingual enhancement have focused on aligning representations in the language-agnostic space (Hu et al., 2024; Berend, 2020; Cao et al., 2020; Karthikeyan et al., 2020; Alaux et al., 2019; Wang et al., 2019) or aligning gradients during optimization (Lee et al., 2022; Wang et al., 2021) to leverage shared features across languages. However, few works in multilingual machine translation have considered language-specific characteristics, primarily to implement routing mechanisms or modular designs to improve performance (Zhao et al., 2024b; Zhang et al., 2021a).
In contrast, our proposed Lens goes a step further that it leverages both language-agnostic and language-specific subspaces to comprehensively enhance multilingual performance both inheriting the theoretical soundness and demonstrating practical utility of our approach.